# The interplay of the parametric instability and MRI in oscillatory shear flows

### v1

@@ 6 Discussion
This study builds upon previous efforts to understand the behavior of instabilities and turbulent stresses in the oscillatory flows established by distorted disks. Notably Torkelsson et al. (2000) also considered the interplay of the MRI and PI in stratified local box simulations. However, they treat the sloshing as a free initial condition, which rapidly decays, compared with our steady forcing setup. They attribute the initial rapid damping to the emergence of the PI, which is then followed by a period of exponential decay as the sloshing interacts with the magnetized turbulence. This allows them to indirectly infer an effective vertical $\alpha\sim 0.006$ which is comparable to the standard horizontal, base MRI stress they find – broadly in support of an isotropic $\alpha$ prescription. However, whilst they measure horizontal stresses which are dominated by the Maxwell component, they are surprised to find vertical stresses which are primarily Reynolds. In fact, this is in keeping with our findings that the strong vertical flow structures set up by the PI survive and are primarily responsible for resisting the sloshing motion. Furthermore, our detailed analysis of the stress versus rate-of-strain relationship advocates for a more viscoelastic behavior which is generally anisotropic between the $\nu_{xz}$ and $\nu_{yz}$ components. Because of the inherent epicylic motions sourced by the MRI, Torkelsson et al. (2000) are fairly restricted to initializing large sloshing motions which stand out above the background. Whilst our fiducial runs also adopt this clear separation, the continual forcing permits us to track the equilibrated sloshing over many dynamical timescales and thereby extract a cleaner signal for small $\psi_{\rm f}$. This yields sloshing responses and viscoelastic amplitudes which vary as a complicated function of $\psi_{\rm f}$, as seen in Fig. 14. Notably, we suggest that the MRI counterintuitively promotes larger sloshing flows as it tends to suppress the vertical hydrodynamic stresses that the PI wants to set up.
@@
Another point of comparison is the purely hydrodynamic, two-dimensional setup considered by Paardekooper & Ogilvie (2019) who followed the growth and saturation of the PI in the warped shearing box formalism of Ogilvie & Latter (2013b). This implementation fully accounts for the warped geometry and stratified disk structure, finding a richness of behavior. Notably the PI can significantly suppress the laminar flows and reduce the efficacy of warp communication. Whilst we ramp up the sloshing from rest, they typically start from the steady laminar flow set by some explicit value of $\alpha$. For larger, $\alpha$ they find that the PI growth is weaker and the effect on the laminar flows is more modest. This seems analogous to the presence of the MRI in our setup which damps the strength of the emergent PI. They also briefly examine the resonant, inviscid limit for which no laminar steady state exists and instead ramp up the sloshing from rest, finding results consistent with their low $\alpha$ runs. The nature of the final sloshing state is found to depend sensitively on which PI radial modal number dominates. Just as we find the emergence of specific elevator modes, they also find the sloshing state is characterized by preferred wavelengths. Note, that these can change in time, since as the sloshing evolves, the modal growth rates vary – hence different modes become preferred and compete for dominance. This can lead to state transitions in the flow and also complex hysteresis effects as the final equilibrium depends on the history of the sloshing. We also observe similar behavior as the strong sloshing initially selects a relatively short wavelength elevator mode (see Fig. 6 panel (a)), before these often merge and are subsumed by longer wavelength channels at late times.
@@
Whilst the work of Paardekooper & Ogilvie (2019) is ideally designed to measure the internal torques relevant to the evolution of warped disks, their two-dimensional simulations perhaps limit the character of the ensuing turbulence and stresses. Whilst our setup is less obviously comparable to the realistic warping scenario, our three-dimensional study, including magnetic fields provides a complementary perspective which proffers more insight into the combined PI-MRI phenomenology. One major caveat is the lack of stratification in our model which makes a direct quantitative link to the full stratified, warping regime tenuous. However, we might be tempted to draw qualitative analogy between regimes when the values of $\psi$ and $\psi_{\rm f}$ predict viscous laminar flows of comparable magnitude. Specifically, for a stratified warped disk, one predicts the sloshing flow induced at $1H$ to be $\sim\psi\,c/\alpha$ (see Eq. (38) of Lodato & Pringle (2007)). Meanwhile, manipulation of (17) yields a maximum sloshing amplitude in our model of $\sim\psi_{\rm f}\,c/(8\pi^{2}\alpha)$. This motivates the correspondence $\psi\sim\psi_{\rm f}/(8\pi^{2})$, for which the equilibrium internal flows and average shear rates are similar across both models. This suggests our fiducial upper limit of $\psi_{\rm f}=0.3$, might relate to lower values of $\psi\sim 0.004$. Even for these small warps, (below typical values for $H/R$) the sloshing motion can become of sonic order owing to the Keplerian resonance. As one progresses to larger warp amplitudes $\psi\sim H/R$, nonlinear effects also come into play wherein the stratified disk can substantially depart from hydrostatic balance. In this case the disk can enter a ‘bouncing’ regime with extreme scale height variations (e.g. Fairbairn & Ogilvie, 2021). This forcing of the disk scale height offers yet another source of modulation to the flow structure which has recently been studied by Held & Ogilvie (2024). This vertical oscillation is also found to be susceptible to another form of the PI, which taps into the free energy associated with the breathing motion, and can generate bending waves. The interplay of the PI modes, driven simultaneously by both the radial sloshing and bouncing motions, merits further investigation.
@@
To this end, in future work we will pursue a stratified extension of this experiment which will allow us to explore some of these neglected effects and map more directly onto the internal flows in realistic warped disks. In particular, upper disk boundaries will confine the elevator flows and presumably modify the sloshing equilibria and associated viscoelastic coefficients.
@@ 7 Conclusion
Motivated by the radial, oscillatory sloshing flows found in warped astrophysical disks, in this paper we have performed the first detailed numerical study examining the interplay of the parametric and magnetorotational instabilities in forced, time-dependent shear flows within a local, unstratified shearing box. In the weak and unforced sloshing regime, we verify the growth of inertial waves in our simulations, in agreement with the predictions of the three-mode coupling theory. We then resonantly drive these sloshing flows and find the resulting equilibrated states across a suite of hydrodynamic and magnetized runs, exploring a range of forcing amplitudes. In both regimes the parametric instability is typically found to dominate the vertical transport of horizontal momentum and temper the sloshing motions. In our hydro runs, the resonantly forced flows initially drive large epicylic oscillations which are unstable to counter-propagating vertical elevator modes, representing the strong sloshing outcome of the parametric instability. Above a critical warping amplitude, the same qualitative behavior emerges in magnetorotationally unstable runs, albeit with slightly increased sloshing amplitude as the MRI turbulence damps the growth of the elevator modes which are then less efficient at resisting the forcing. Below this critical amplitude, the elevator modes are completely suppressed and the sloshing motion is mediated primarily by the MRI. A detailed comparison of the stress and rate-of-strain profiles suggests that the vertical stresses are more accurately described by anisotropic, viscoelastic coefficients as opposed to the simplistic $\alpha$ description. Nonetheless, over the range of forcings considered, the extracted vertical viscoelastic coefficients are of similar amplitude to the horizontal $\alpha$ viscosity associated with the standard MRI, despite their physical characters being emphatically distinct. This work paves the way for future investigation into magnetized warps, where we will extend our setup to include stratification, allowing a more direct connection with realistic distorted disks.
@@ acknowledgments
The authors would like to thank Gordon Ogilvie, Geoffroy Lesur and Martin Pessah for helpful conversations during the preparation of this manuscript. This research was supported by the W. M. Keck Foundation Fund and the IAS Fund for Memberships in the Natural Sciences (CWF).
@@ Appendix A Parametric Instability Theory
In order to benchmark our numerical setup and its capacity to simulate oscillatory, shear flows, it is instructive to develop an analytical theory for the dynamics of weakly-nonlinear sloshing. Indeed, it has long been recognized that time-dependent shear flows are susceptible to a three mode-coupling, parametric instability (Gammie et al., 2000) which will be developed in detail for our unique setup in this section. Here, we will broadly follow the analysis expounded in Ogilvie & Latter (2013b) and Fairbairn & Ogilvie (2023) which instead dealt with isothermal and polytropic, stratified disks.
@@ A.1 Governing equations
where we have introduced the material derivative, $D\equiv\partial_{t}+\mathbf{u}\cdot\nabla$ and our thermodynamic quantities are replaced with the pseudo-enthalpy variable $h=c^{2}\ln\rho$. As described in section 2, these equations readily admit the solution,
@@
We now Fourier decompose the perturbations according to $X^{\prime}=\sum_{n=-\infty}^{\infty}\tilde{X}_{n}\exp{\left[i(k_{x}x+2\pi nz/L_{z})\right]}$. Inserting this modal ansatz into equations (A8)-(A11), we exploit the orthogonality of complex exponentials to project onto evolutionary equations for each vertical order $n$. Upon non-dimensionalising the problem we arrive at
@@
$\partial_{\tau}h_{n}-(K_{x}S/2)\cos(\tau)\left(h_{n+1}-h_{n-1}\right)+iK_{x}u_{n}+iK_{\rm b}nw_{n}=0.$ (A15)
@@ A.2.1 Leading order: $\mathcal{O}(S^{0})$
$\mathbf{L}_{n}=\begin{pmatrix}\partial_{0}&-2&0&iK_{x}\\ 1/2&\partial_{0}&0&0\\ 0&0&\partial_{0}&iK_{b}n\\ iK_{x}&0&iK_{b}n&\partial_{0}\end{pmatrix}.$ (A17)
@@
$(-\omega_{0}^{2}+K_{b}^{2}n^{2})(-\omega_{0}^{2}+1)-K_{x}^{2}\omega_{0}^{2}=0.$ (A18)
This describes a family of frequency curves as functions of radial wavenumber, each branch labeled by the discrete vertical mode number $n$. For visualization, these inertial branches are plotted in Fig. 2 as different colors. The blue branch denotes the non-oscillatory $n=0$ mode with eigenvectors satisfying $v_{n}^{(0)}=iK_{x}h_{n}^{(0)}/2$, $u_{n}^{(0)}=0$. This steady-state mode corresponds to an adjusted geostrophic balance between the radial pressure (enthalpy) gradients and the Coriolis force, induced through the perturbed azimuthal velocity. Meanwhile, the vertical velocity can flow freely with arbitrary $w_{n}^{(0)}$. In fact, this linear mode is also a nonlinear solution to the full equations (A1)-(A4). Although it does not participate in the weakly nonlinear parametric instability, it is clearly emergent in our strong sloshing numerical experiments (see section 4.2). The lowest frequency oscillatory branch labels the $n=1$ vertical mode, which incrementally increases toward higher $n$ as one jumps to higher frequency branches at a given $K_{x}$. The eigenmodes associated with the $n\geq 1$ branches are described by
$\mathbf{V}_{n}^{(0)}(\omega;K_{x})=\begin{bmatrix}&i\omega(\omega^{2}-K_{b}^{2}n^{2})\\ &\frac{1}{2}(\omega^{2}-K_{b}^{2}n^{2})\\ &iK_{x}K_{b}n\omega\\ &iK_{x}\omega^{2}\end{bmatrix}e^{-i\omega\tau_{0}}.$ (A19)
@@
$\mathcal{L}_{n}[h_{n}^{(0)}]=\left[(\partial_{0}^{2}+K_{b}^{2}n^{2})(\partial_{0}^{2}+1)-K_{x}^{2}\partial_{0}^{2}\right]h_{n}^{(0)}=0.$ (A20)
@@ A.2.2 First order: $\mathcal{O}(S^{1})$
where the right hand side forcing vector is given by
$\mathbf{F}_{n}^{(1)}=-\partial_{1}\mathbf{V}_{n}^{(0)}+\begin{bmatrix}&\frac{1}{2}K_{x}\cos{\tau_{0}}(u_{n+1}^{(0)}-u_{n-1}^{(0)})-\frac{1}{2}K_{b}\cos\tau_{0}(w_{n+1}^{(0)}+w_{n-1}^{(0)})\\ &\frac{1}{2}K_{x}\cos{\tau_{0}}(v_{n+1}^{(0)}-v_{n-1}^{(0)})+\frac{1}{4}K_{b}\sin\tau_{0}(w_{n+1}^{(0)}+w_{n-1}^{(0)})\\ &\frac{1}{2}K_{x}\cos\tau_{0}(w_{n+1}^{(0)}-w_{n-1}^{(0)})\\ &\frac{1}{2}K_{x}\cos\tau_{0}(h_{n+1}^{(0)}+h_{n-1}^{(0)})\end{bmatrix}.$ (A22)
The inertial wave operator $\mathbf{L}_{n}$ is inherited from leading order and now acts on the first-order perturbations. The wave modes from zeroth order in $S$ now manifest as forcing contributions on the right hand side. Notably, this includes ‘off-diagonal’ coupling terms since the product of the $n=1$ sloshing structure with $n^{\rm th}$ order waves, facilitates communication with the neighboring $n\pm 1$ modes.
Since the singular left hand side operator obviously admits inertial waves as non-trivial solutions, the forcing vector on the right hand side must be orthogonal to these solutions in some sense. More specifically, consider combining the rows of equation (A21) into differential operator form. This yields an equation of the form
@@
Since the left hand side is simply the inertial wave operator from zeroth order for the $(K_{x},n)$ mode, the right hand side must contain no terms proportional to $e^{-i\omega_{0}\tau_{0}}$ which would otherwise resonantly force a singular, nonphysical response. If the forcing vector $\mathbf{F}_{n}^{(1)}$ contains such resonant contributions, which we denote $\mathbf{F}_{n,\omega_{0}}^{(1)}=[A,B,C,D]^{\mathsf{T}}$, then in order for these to cancel upon combination under the row operations, we must impose the solvability condition
$K_{x}\omega_{0}(\omega_{0}A+2iB)+(\omega_{0}^{2}-1)(K_{b}nC+\omega_{0}D)=0.$ (A24)
This solvability condition will lead to a coupling between the wave modes found at zeroth order. Specifically, the relevant modal combination should involve a three-mode resonance, where the parent sloshing wave interacts with two inertial daughter modes, with dimensionless angular frequencies denoted $\omega_{1}$ and $\omega_{2}$ and vertical modal numbers labeled $n_{1}$ and $n_{2}$. Since the sloshing is horizontally homogeneous, the inertial modes should share the same $K_{x}$. Meanwhile, in order to overlap with the vertical and temporal structure of the sloshing, we require $\omega_{2}-\omega_{1}=1$ and $n_{2}-n_{1}=1$. Such resonant couplings are shown by the connective lines between neighboring branches of the dispersion relation curves in Fig. 2. Constructing a modal sum that satisfies these conditions yields
@@
This modal sum can be inserted into equation (A22) for $\mathbf{F}_{n}^{(1)}$ and $\mathbf{F}_{n+1}^{(1)}$. In each case, we extract the terms proportional to $e^{i\omega_{0}\tau_{0}}$ and $e^{i(\omega_{0}+1)\tau_{0}}$ which would resonate with the left hand side operator, giving
$\mathbf{F}_{n,\omega_{0}}^{(1)}=-\partial_{1}\mathbf{U}_{n}^{(0)}+\frac{1}{4}\begin{bmatrix}K_{x}U_{n+1}^{(0)}-K_{b}W_{n+1}^{(0)}\\ K_{x}V_{n+1}^{(0)}-\frac{i}{2}K_{b}W_{n+1}^{(0)}\\ K_{x}W_{n+1}^{(0)}\\ K_{x}h_{n+1}^{(0)}\\ \end{bmatrix}$ (A29)
@@
$\mathbf{F}_{n+1,\omega_{0}+1}^{(1)}=-\partial_{1}\mathbf{U}_{n+1}^{(0)}-\frac{1}{4}\begin{bmatrix}K_{x}U_{n}^{(0)}+K_{b}W_{n}^{(0)}\\ K_{x}V_{n}^{(0)}-\frac{i}{2}K_{b}W_{n}^{(0)}\\ K_{x}W_{n}^{(0)}\\ K_{x}h_{n}^{(0)}\\ \end{bmatrix}.$ (A30)
@@
$C_{1}=\frac{K_{x}(2\omega_{0}+1)\left(-K_{b}^{2}\omega_{0}^{2}-K_{b}^{2}n^{2}-K_{b}^{2}n\omega_{0}^{2}-K_{b}^{2}n+\omega_{0}^{4}+2\omega_{0}^{3}+\omega_{0}^{2}\right)}{8\omega_{0}\left(\omega_{0}^{4}-K_{b}^{2}n^{2}\right)}$ (A33)
$C_{2}=\frac{K_{x}(2\omega_{0}+1)\left(-K_{b}^{2}n^{2}+K_{b}^{2}n\omega_{0}(\omega_{0}+2)+\omega_{0}^{2}(\omega_{0}+1)^{2}\right)}{8(\omega_{0}+1)\left(K_{b}^{2}(n+1)^{2}-(\omega_{0}+1)^{4}\right)}.$ (A34)

### v2

@@ 6 Discussion
This study builds upon previous efforts to understand the behavior of instabilities and turbulent stresses in the oscillatory flows established by distorted disks. Notably, Torkelsson et al. (2000) also considered the interplay of the MRI and PI in stratified local box simulations. However, they treat the sloshing as a free initial condition, which rapidly decays, compared with our steady forcing setup. They attribute the initial rapid damping to the emergence of the PI, which is then followed by a period of exponential decay as the sloshing interacts with the magnetized turbulence. This allows them to indirectly infer an effective vertical $\alpha\sim 0.006$, which is comparable to the standard horizontal, base MRI stress they find – broadly in support of an isotropic $\alpha$ prescription. However, whilst they measure horizontal stresses, which are dominated by the Maxwell component, they are surprised to find vertical stresses, which are primarily Reynolds. In fact, this is in keeping with our findings that the strong vertical flow structures set up by the PI survive and are primarily responsible for resisting the sloshing motion. Furthermore, our detailed analysis of the stress versus rate of strain relationship advocates for a more viscoelastic behavior, which is generally anisotropic between the $\nu_{xz}$ and $\nu_{yz}$ components. Because of the inherent epicylic motions sourced by the MRI, Torkelsson et al. (2000) are fairly restricted to initializing large sloshing motions which stand out above the background. Whilst our fiducial runs also adopt this clear separation, the continual forcing permits us to track the equilibrated sloshing over many dynamical timescales and thereby extract a cleaner signal for small $\psi_{\rm f}$. This yields sloshing responses and viscoelastic amplitudes which vary as a complicated function of $\psi_{\rm f}$, as seen in Fig. 14. Notably, we suggest that the MRI counterintuitively promotes larger sloshing flows as it tends to suppress the vertical hydrodynamic stresses that the PI wants to set up.
@@
Another point of comparison is the purely hydrodynamic, two-dimensional setup considered by Paardekooper & Ogilvie (2019), who followed the growth and saturation of the PI in the warped shearing box formalism of Ogilvie & Latter (2013b). This implementation fully accounts for the warped geometry and stratified disk structure, finding a richness of behavior. Notably, the PI can significantly suppress the laminar flows and reduce the efficacy of warp communication. Whilst we ramp up the sloshing from rest, they typically start from the steady laminar flow set by some explicit value of $\alpha$. For larger $\alpha$, they find that the PI growth is weaker and the effect on the laminar flows is more modest. This seems analogous to the presence of the MRI in our setup, which damps the strength of the emergent PI. They also briefly examine the resonant, inviscid limit for which no laminar steady state exists and instead ramp up the sloshing from rest, finding results consistent with their low $\alpha$ runs. The nature of the final sloshing state is found to depend sensitively on which PI radial modal number dominates. Just as we find the emergence of specific elevator modes, they also find that the sloshing state is characterized by preferred wavelengths. Note that these can change in time, since as the sloshing evolves, the modal growth rates vary – hence different modes become preferred and compete for dominance. This can lead to state transitions in the flow and also complex hysteresis effects as the final equilibrium depends on the history of the sloshing. We also observe similar behavior as the strong sloshing initially selects a relatively short wavelength elevator mode (see Fig. 6 panel (a)), before these often merge and are subsumed by longer wavelength channels at late times.
@@
Whilst the work of Paardekooper & Ogilvie (2019) is ideally designed to measure the internal torques relevant to the evolution of warped disks; their two-dimensional simulations perhaps limit the character of the ensuing turbulence and stresses. Whilst our setup is less obviously comparable to the realistic warping scenario, our three-dimensional study, including magnetic fields, provides a complementary perspective that proffers more insight into the combined PI-MRI phenomenology. One major caveat is the lack of stratification in our model, which makes a direct quantitative link to the full stratified, warping regime tenuous. However, we might be tempted to draw a qualitative analogy between regimes when the values of $\psi$ and $\psi_{\rm f}$ predict viscous laminar flows of comparable magnitude. In particular, for a stratified warped disk, one predicts the sloshing flow induced at $1H$ to be $\sim\psi\,c/\alpha$ (see Eq. (38) of Lodato & Pringle (2007)). Meanwhile, manipulation of (17) yields a maximum sloshing amplitude in our model of $\sim\psi_{\rm f}\,c/(8\pi^{2}\alpha)$. This motivates the correspondence $\psi\sim\psi_{\rm f}/(8\pi^{2})$, for which the equilibrium internal flows and average shear rates are similar across both models. This suggests our fiducial upper limit of $\psi_{\rm f}=0.3$, might relate to lower values of $\psi\sim 0.004$. Even for these small warps (below typical values for $H/R$), the sloshing motion can become of sonic order owing to the Keplerian resonance. As one progresses to larger warp amplitudes $\psi\sim H/R$, nonlinear effects also come into play wherein the stratified disk can substantially depart from hydrostatic balance. In this case, the disk can enter a ‘bouncing’ regime with extreme scale height variations (e.g. Fairbairn & Ogilvie, 2021). This forcing of the disk scale height offers yet another source of modulation to the flow structure, which has recently been studied by Held & Ogilvie (2024). This vertical oscillation is also found to be susceptible to another form of the PI, which taps into the free energy associated with the breathing motion and can generate bending waves. The interplay of the PI modes, driven simultaneously by both the radial sloshing and bouncing motions, merits further investigation.
@@
To this end, in future work, we will pursue a stratified extension of this experiment, which will allow us to explore some of these neglected effects and map more directly onto the internal flows in realistic warped disks. In particular, upper disk boundaries will confine the elevator flows and presumably modify the sloshing equilibria and associated viscoelastic coefficients.
@@ 7 Conclusion
Motivated by the radial, oscillatory sloshing flows found in warped astrophysical disks, in this paper, we have performed the first detailed numerical study examining the interplay of the parametric and magnetorotational instabilities in forced, time-dependent shear flows within a local, unstratified shearing box. In the weak and unforced sloshing regime, we verify the growth of inertial waves in our simulations, in agreement with the predictions of the three-mode coupling theory. We then resonantly drive these sloshing flows and find the resulting equilibrated states across a suite of hydrodynamic and magnetized runs, exploring a range of forcing amplitudes. In both regimes, the parametric instability is typically found to dominate the vertical transport of horizontal momentum and temper the sloshing motions. In our hydrodynamic runs, the resonantly forced flows initially drive large epicylic oscillations, which are unstable to counter-propagating vertical elevator modes, representing the strong sloshing outcome of the parametric instability. Above a critical warping amplitude, the same qualitative behavior emerges in magnetorotationally unstable runs, albeit with slightly increased sloshing amplitude as the MRI turbulence damps the growth of the elevator modes, which are then less efficient at resisting the forcing. Below this critical amplitude, the elevator modes are completely suppressed, and the sloshing motion is mediated primarily by the MRI. A detailed comparison of the stress and rate of strain profiles suggests that the vertical stresses are more accurately described by anisotropic, viscoelastic coefficients as opposed to the simplistic $\alpha$ description. Nonetheless, over the range of forcings considered, the extracted vertical viscoelastic coefficients are of similar amplitude to the horizontal $\alpha$ viscosity associated with the standard MRI, despite their physical characters being emphatically distinct. This work paves the way for future investigation into magnetized warps, where we will extend our setup to include stratification, allowing for a more direct connection with realistic distorted disks.
@@ Acknowledgments
The authors thank the referee for a detailed and thought-provoking report. We would also like to thank Gordon Ogilvie, Geoffroy Lesur, and Martin Pessah for helpful conversations during the preparation of this manuscript. This research was supported by the W. M. Keck Foundation Fund and the IAS Fund for Memberships in the Natural Sciences (CWF).
@@ Appendix A Parametric Instability Theory
In order to benchmark our numerical setup and its capacity to simulate oscillatory, shear flows, it is instructive to develop an analytical theory for the dynamics of weakly nonlinear sloshing. Indeed, it has long been recognized that time-dependent shear flows are susceptible to a three-mode-coupling, parametric instability (Gammie et al., 2000), which will be developed in detail for our unique setup in this section. Here, we will broadly follow the analysis expounded in Ogilvie & Latter (2013b) and Fairbairn & Ogilvie (2023), which instead dealt with isothermal and polytropic, stratified disks.
@@ A.1 Governing Equations
where we have introduced the material derivative, $D\equiv\partial_{t}+\mathbf{u}\cdot\nabla$, and our thermodynamic quantities are replaced with the pseudo-enthalpy variable $h=c^{2}\ln\rho$. As described in Section 2, these equations readily admit the solution
@@
We now Fourier decompose the perturbations according to $X^{\prime}=\sum_{n=-\infty}^{\infty}\tilde{X}_{n}\exp{\left[i(k_{x}x+2\pi nz/L_{z})\right]}$. Inserting this modal ansatz into equations (A8)-(A11), we exploit the orthogonality of complex exponentials to project onto evolutionary equations for each vertical order $n$. Upon nondimensionalizing the problem we arrive at
@@
$\partial_{\tau}h_{n}-(K_{x}S/2)\cos(\tau)\left(h_{n+1}-h_{n-1}\right)+iK_{x}u_{n}+iK_{\rm b}nw_{n}=0,$ (A15)
@@ A.2.1 Leading Order: $\mathcal{O}(S^{0})$
$\mathbf{L}_{n}=\begin{pmatrix}\partial_{0}&-2&0&iK_{x}\\ 1/2&\partial_{0}&0&0\\ 0&0&\partial_{0}&iK_{\rm b}n\\ iK_{x}&0&iK_{\rm b}n&\partial_{0}\end{pmatrix}$. (A17)
@@
$(-\omega_{0}^{2}+K_{\rm b}^{2}n^{2})(-\omega_{0}^{2}+1)-K_{x}^{2}\omega_{0}^{2}=0.$ (A18)
This describes a family of frequency curves as functions of the radial wavenumber, each branch labeled by the discrete vertical mode number $n$. For visualization, these inertial branches are plotted in Fig. 2 as different colors. The blue branch denotes the nonoscillatory $n=0$ mode with eigenvectors satisfying $v_{n}^{(0)}=iK_{x}h_{n}^{(0)}/2$, $u_{n}^{(0)}=0$. This steady-state mode corresponds to an adjusted geostrophic balance between the radial pressure (enthalpy) gradients and the Coriolis force, induced through the perturbed azimuthal velocity. Meanwhile, the vertical velocity can flow freely with arbitrary $w_{n}^{(0)}$. In fact, this linear mode is also a nonlinear solution to the full equations (A1)-(A4). Although it does not participate in the weakly nonlinear parametric instability, it is clearly emergent in our strong sloshing numerical experiments (see Section 4.2). The lowest frequency oscillatory branch labels the $n=1$ vertical mode, which incrementally increases toward higher $n$ as one jumps to higher frequency branches at a given $K_{x}$. The eigenmodes associated with the $n\geq 1$ branches are described by
$\mathbf{V}_{n}^{(0)}(\omega;K_{x})=\begin{bmatrix}&i\omega(\omega^{2}-K_{\rm b}^{2}n^{2})\\ &\frac{1}{2}(\omega^{2}-K_{\rm b}^{2}n^{2})\\ &iK_{x}K_{\rm b}n\omega\\ &iK_{x}\omega^{2}\end{bmatrix}e^{-i\omega\tau_{0}}.$ (A19)
@@
$\mathcal{L}_{n}[h_{n}^{(0)}]=\left[(\partial_{0}^{2}+K_{\rm b}^{2}n^{2})(\partial_{0}^{2}+1)-K_{x}^{2}\partial_{0}^{2}\right]h_{n}^{(0)}=0.$ (A20)
@@ A.2.2 First Order: $\mathcal{O}(S^{1})$
where the right-hand side forcing vector is given by
$\mathbf{F}_{n}^{(1)}=-\partial_{1}\mathbf{V}_{n}^{(0)}+\begin{bmatrix}&\frac{1}{2}K_{x}\cos{\tau_{0}}(u_{n+1}^{(0)}-u_{n-1}^{(0)})-\frac{1}{2}K_{\rm b}\cos\tau_{0}(w_{n+1}^{(0)}+w_{n-1}^{(0)})\\ &\frac{1}{2}K_{x}\cos{\tau_{0}}(v_{n+1}^{(0)}-v_{n-1}^{(0)})+\frac{1}{4}K_{\rm b}\sin\tau_{0}(w_{n+1}^{(0)}+w_{n-1}^{(0)})\\ &\frac{1}{2}K_{x}\cos\tau_{0}(w_{n+1}^{(0)}-w_{n-1}^{(0)})\\ &\frac{1}{2}K_{x}\cos\tau_{0}(h_{n+1}^{(0)}+h_{n-1}^{(0)})\end{bmatrix}$. (A22)
The inertial wave operator $\mathbf{L}_{n}$ is inherited from leading order and now acts on the first-order perturbations. The wave modes from zeroth order in $S$ now manifest as forcing contributions on the right-hand side. Notably, this includes ‘off-diagonal’ coupling terms since the product of the $n=1$ sloshing structure with $n^{\rm th}$ order waves, facilitates communication with the neighboring $n\pm 1$ modes.
Since the singular left-hand side operator obviously admits inertial waves as nontrivial solutions, the forcing vector on the right-hand side must be orthogonal to these solutions in some sense. More specifically, consider combining the rows of equation (A21) into differential operator form. This yields an equation of the form
@@
Since the left-hand side is simply the inertial wave operator from zeroth order for the $(K_{x},n)$ mode, the right-hand side must contain no terms proportional to $e^{-i\omega_{0}\tau_{0}}$, which would otherwise resonantly force a singular, nonphysical response. If the forcing vector $\mathbf{F}_{n}^{(1)}$ contains such resonant contributions, which we denote $\mathbf{F}_{n,\omega_{0}}^{(1)}=[A,B,C,D]^{\mathsf{T}}$, then in order for these to cancel upon combination under the row operations, we must impose the solvability condition
$K_{x}\omega_{0}(\omega_{0}A+2iB)+(\omega_{0}^{2}-1)(K_{\rm b}nC+\omega_{0}D)=0.$ (A24)
This solvability condition will lead to a coupling between the wave modes found at zeroth order. In particular, the relevant modal combination should involve a three-mode resonance, where the parent sloshing wave interacts with two inertial daughter modes, with dimensionless angular frequencies denoted $\omega_{1}$ and $\omega_{2}$ and vertical modal numbers labeled $n_{1}$ and $n_{2}$. Since the sloshing is horizontally homogeneous, the inertial modes should share the same $K_{x}$. Meanwhile, in order to overlap with the vertical and temporal structure of the sloshing, we require $\omega_{2}-\omega_{1}=1$ and $n_{2}-n_{1}=1$. Such resonant couplings are shown by the connective lines between neighboring branches of the dispersion relation curves in Fig. 2. Constructing a modal sum that satisfies these conditions yields
@@
This modal sum can be inserted into equation (A22) for $\mathbf{F}_{n}^{(1)}$ and $\mathbf{F}_{n+1}^{(1)}$. In each case, we extract the terms proportional to $e^{i\omega_{0}\tau_{0}}$ and $e^{i(\omega_{0}+1)\tau_{0}}$, which would resonate with the left-hand side operator, giving
$\mathbf{F}_{n,\omega_{0}}^{(1)}=-\partial_{1}\mathbf{U}_{n}^{(0)}+\frac{1}{4}\begin{bmatrix}K_{x}U_{n+1}^{(0)}-K_{\rm b}W_{n+1}^{(0)}\\ K_{x}V_{n+1}^{(0)}-\frac{i}{2}K_{\rm b}W_{n+1}^{(0)}\\ K_{x}W_{n+1}^{(0)}\\ K_{x}h_{n+1}^{(0)}\\ \end{bmatrix}$ (A28)
@@
$\mathbf{F}_{n+1,\omega_{0}+1}^{(1)}=-\partial_{1}\mathbf{U}_{n+1}^{(0)}-\frac{1}{4}\begin{bmatrix}K_{x}U_{n}^{(0)}+K_{\rm b}W_{n}^{(0)}\\ K_{x}V_{n}^{(0)}-\frac{i}{2}K_{\rm b}W_{n}^{(0)}\\ K_{x}W_{n}^{(0)}\\ K_{x}h_{n}^{(0)}\\ \end{bmatrix}$. (A29)
@@
$C_{1}=\frac{K_{x}(2\omega_{0}+1)\left(-K_{\rm b}^{2}\omega_{0}^{2}-K_{\rm b}^{2}n^{2}-K_{\rm b}^{2}n\omega_{0}^{2}-K_{\rm b}^{2}n+\omega_{0}^{4}+2\omega_{0}^{3}+\omega_{0}^{2}\right)}{8\omega_{0}\left(\omega_{0}^{4}-K_{\rm b}^{2}n^{2}\right)}$ (A32)
$C_{2}=\frac{K_{x}(2\omega_{0}+1)\left(-K_{\rm b}^{2}n^{2}+K_{\rm b}^{2}n\omega_{0}(\omega_{0}+2)+\omega_{0}^{2}(\omega_{0}+1)^{2}\right)}{8(\omega_{0}+1)\left(K_{\rm b}^{2}(n+1)^{2}-(\omega_{0}+1)^{4}\right)}.$ (A33)